What is the step by step way of solving -190=8+6(-1+4n)

Answer by mxgirl22(39)



-190=8+6(-1+4n)
-190=8-6+24n
-190=2+24n
-190-2=2-2+24n
-192=24n
-192/24=24n/24
-8=n